# supplement

### v1

@@ 5.3 Optimized mode: distributing the sampling budget across prompts to reduce
Having bounded $\gamma$, our next objective is to minimize the average weight $\bar{w}=\frac{1}{|{\mathcal{I}_{2}}|}\sum_{i\in\mathcal{I}_{2}}\frac{1}{\pi_{i}}$ in order to reduce the variance of $\hat{\alpha}(\tau)$. Recall the budget constraint in (1) and observe that $\mathbb{E}[\sum_{i\in\mathcal{I}_{2}}C_{i}|\{X_{i}\}_{i\in{\mathcal{I}_{2}}}]=\sum_{i\in\mathcal{I}_{2}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\,\pi_{i}$ by construction. In turn, we can minimize $\bar{w}$ subject to the budget constraint by solving the following convex optimization problem:
$\pi^{*}=\underset{{\pi\in[0,1]^{|{\mathcal{I}_{2}}|}}}{\rm{argmin}}\;\frac{1}{|{\mathcal{I}_{2}}|}\sum_{i\in\mathcal{I}_{2}}\frac{1}{\pi_{i}}\quad\text{s.t.}\quad\sum_{i\in\mathcal{I}_{2}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\,\pi_{i}\!\leq\!B.$ (8)
@@
In Appendix B.2, we introduce Algorithm 4, which efficiently solves the above optimization problem. Proposition B.2 ensures this algorithm returns a unique and strictly positive solution $\pi^{*}$ that satisfies the budget constraint (1). Importantly, because the solution $\pi^{*}$ depends on the entire set of calibration prompts $\{X_{i}\}_{i\in{\mathcal{I}_{2}}}$, the resulting weights inherit this dependence. To make this explicit, we denote the weight associated with the $i$-th example as $w(\{X_{j}\}_{j\in\mathcal{I}_{2}},i)=1/\pi_{i}^{*}$. Notably, these weights are also guaranteed to be upper-bounded by the same expression for $\gamma$ as in the Trimmed method.
@@
Suppose that $\max_{i\in{\mathcal{I}_{2}}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\leq M$, then, the weights $w(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)$ induced by solving (8) are upper bounded by $\gamma=\max\left({|{\mathcal{I}_{2}}|\cdot M}/{B},1\right).$
All proofs are deferred to Appendix D. The above proposition shows us that the solution $\pi^{*}$ inherits the advantages of the Trimmed method and further improves upon it by achieving lower variance, as it minimizes $\bar{w}$.
The Optimized calibration scheme is outlined in Algorithm 2, and its validity is stated next. The following result builds on the theory of [15] and extends it to accommodate the dependency structure of the weights induced by $\pi^{*}$.
Theorem 5.2 (Validity of the prompt-adaptive methods—Basic, Trimmed, and Optimized).
Fix a tolerance level $\delta\in\left(0,1\right)$ and a miscoverage level $\tau\in\left(0,1\right)$. Suppose that $\{(X_{i},T_{i})\}_{i\in{\mathcal{I}_{2}}}$ and $(X_{\textup{test}},T_{\textup{test}})$ are drawn i.i.d. We further suppose that $\hat{q}_{\tau}(x)$ is non-decreasing and continuous in $\tau$, and that there exists a constant ${\gamma}>0$ such that ${w}(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)\leq{\gamma}$ for all $i\in{\mathcal{I}_{2}}$ and for $P_{X}$-almost all $\{X_{j}\}_{j\in{\mathcal{I}_{2}}}$. Then, with probability at least $1-\delta$ over the draws of $\mathcal{D}$, the LPB $\hat{L}$ produced by Algorithm 2 attains a valid coverage rate:
$\mathbb{P}\left[T_{\textup{test}}\geq\hat{L}(X_{\textup{test}})|\mathcal{D}\right]\geq 1-\alpha-\sqrt{\frac{2\gamma^{2}+5}{|{\mathcal{I}_{2}}|}\cdot\log\left(\frac{1}{\delta}\right)}.$ (9)
@@ 6 Experiments
We evaluate our methods on both synthetic (Section 6.1) and real (Section 6.2) datasets. For a full description of how we train models to estimate the conditional quantiles of $T$ given $X$, along with additional experiments, implementation details, and runtime considerations for our calibration methods, see Appendix E.
@@ 6.1 Synthetic data experiments
We generate a dataset that simulates a high-risk setting with predominantly prompts that can yield unsafe outcomes. To this end, we generate $n=100{,}000$ pairs $(X_{i},p_{i})$, where $p_{i}$ is the true probability of sampling an unsafe $Y_{i}$ for $X_{i}\in\mathbb{R}^{d}$, with $d=10$. Specifically, $X_{i}\mid p_{i},$ follows a normal distribution, defined in Appendix E.1. The unsafe probability $p$ is chosen to simulate a dataset of “suspicious” prompts—90% of which have a high probability for unsafe generation (a log-space between $10^{-3}$ and $10^{-4}$), with the remaining 10% of prompts having a low probability for such an event (a log-space between $10^{-5}$ and $10^{-6}$). Then, we draw $T_{i}\sim\mathrm{Geom}(p_{i})$. We randomly partition the data into training (45%), calibration (45%), and test (10%) sets. For each example in the training set, we generate $500$ outputs. We then use this training set to fit the quantile estimators $\hat{q}_{\tau}(x)$.
@@
We compare five calibration methods: (1) Uncalibrated baseline, being the raw quantile estimator $\hat{q}_{\tau}(X)$ at quantile level $\tau=\alpha$; (2) the Naive calibration method, which is a slight variant of the method from Section 4. We found the original Naive approach to be highly unstable, generating trivial LPBs. This is because the number of points for which $\hat{q}(X_{i})\leq C_{i}$ is vanishingly small for high quantile levels. As such, to avoid numerical instabilities, we restrict the $\tau$’s search space in Algorithm 1 to $\mathcal{T}\cap[0,{\tau_{\textup{prior}}}]$, as in the prompt-adaptive methods. We also include (3) the Basic prompt-adaptive method from Section 5.1; (4) its Trimmed version from Section 5.2; and (5) our flagship Optimized approach from Section 5.3. For these methods, we set the prior quantile at $\tau_{\mathrm{prior}}=10^{-1/4}$. For the Trimmed and Optimized methods we choose $M$ so that $\gamma=\max\left({|{\mathcal{I}_{2}}|\cdot M}/{B},1\right)=10$.
@@
Recall that the calibration procedures are inherently random due to the random sampling of $C_{i}$ and the random generations obtained by the model $\mathcal{G}(x)$. Ideally, the randomness induced by the calibration procedure should have a small effect on our bounds. To quantify this, we fix the train/calibration/test split, and run each calibration method for $J=20$ random draws of censoring and time-to-unsafe-sampling times. We index each calibration run by $j$, and denote the resulting lower prediction bound by $\hat{q}_{\hat{\tau}}^{(j)}(X_{i})$. Then, for the $j$-th run, we compute:
$\mathrm{AvgCoverage}^{(j)}=\frac{1}{|\mathcal{I}_{\mathrm{test}}|}\sum_{i\in\mathcal{I}_{\mathrm{test}}}$$\mathbb{P}\{T_{i}\leq\hat{q}_{\hat{\tau}}^{(j)}(X_{i})|X_{i}\},\mathrm{AvgLPB}^{(j)}=\frac{1}{|\mathcal{I}_{\mathrm{test}}|}\sum_{i\in\mathcal{I}_{\mathrm{test}}}\hat{q}_{\hat{\tau}}^{(j)}(X_{i}),$
$\mathrm{AvgBudget}^{(j)}=\frac{1}{|\mathcal{I}_{\mathrm{test}}|}\sum_{i\in\mathcal{I}_{\mathrm{test}}}C_{i}^{(j)}.$
@@
Here, $\mathcal{I}_{\mathrm{test}}$ are the indexes of the test points. We report the average and standard deviation of the above quantities over $J$ runs.
@@
Results. We study the effect of the average budget per prompt ($B/|{\mathcal{I}_{2}}|$) on the actual coverage and mean LPB obtained by each method. Following Figure 1, we can see that the Uncalibrated method does not yield valid coverage. The Naive method severely undercovers $T$ at the low budget regime, and overcovers $T$ for larger sampling budgets. The Basic method attains more regulated coverage around the desired $90\%$ level (except the low-budget regime), but has high variance. By contrast, the Trimmed method attains nearly $90\%$ coverage across all budget constraints. Yet, our flagship Optimized calibration method attains tight 90% coverage with minimal variance compared to all other methods. The middle panel highlights that both the Naive and Optimized approaches exhaust the entire budget per sample, however, the latter is tightly regulated around the desired coverage—better utilizing the sampling budget. The right panel demonstrates that the Trimmed and Optimized methods produce stable LPBs with improved power as the average budget per prompt increases. The increased budget allows us to set higher quantile trim values $M$ while enforcing low $\gamma$. Observe that the two adaptive methods get closer to the Basic method as the budget increases. Observe also that the latter produces higher LPBs but with high coverage variability.
Additional results, including varying maximal weights, $\alpha$ levels, and calibration/test splits, are in Appendix E.1.
@@ 6.2 Real data experiments
We now turn to demonstrate the effectiveness of our Optimized calibration method on real-world data. We perform experiments using the RealToxicityPrompts dataset [11], which contains 99,442 sentences used to assess model toxicity. Each sentence in the dataset is split into two halves. The first half of each sentence serves as the input prompt $X_{i}$ to an LLM (we use Llama 3.2 1B [40, 14]), whose task is to complete the sentence. Since we do not have access to human annotations, we use the Detoxify-original model [18], with a $0.5$ toxicity score threshold to classify each generated sentence as safe (non-toxic) or unsafe (toxic).
We randomly split the dataset into training (50%), validation (10%), calibration (20%), and test (20%) sets. On the training set, we sample $N=500$ responses per prompt and fit a quantile estimator $\hat{q}_{\tau}(x)$ as detailed in Appendix E. We set a miscoverage level $\alpha=0.1$ and compare three calibration methods described in Section 6.1: (1) the Uncalibrated baseline, (2) the Naive method, and (3) the Optimized method. We set $\tau_{\mathrm{prior}}=10^{-1/4}$. For any given budget $B$, we set $M$ such that $\gamma=2$.
In all experiments, we report the empirical mean of the LPBs over the test set. We also estimate the empirical miscoverage by drawing $\min\bigl{(}\hat{L}(X_{i}),2400\bigr{)}$ samples for each prompt. We cap the maximal LLM generations on the test set at 2,400 samplings per prompt due to computational constraints. Since the Optimized method trims the quantiles at $M<2400$ across all evaluated budget levels, by drawing exactly $\hat{L}(X_{i})$ samples (with $\hat{L}(X_{i})\leq M$) we can compute the miscoverage rate directly. Since $\hat{L}(X_{i})$ is unbounded for the Uncalibrated and Fixed methods, we can only report a lower bound on the empirical coverage rate using the 2,400 generations; see also [15]. In turn, the empirical lower and upper bounds are given by $\frac{1}{|\mathcal{I}_{\mathrm{test}}|}\sum_{i\in\mathcal{I}_{\mathrm{test}}}\mathbb{I}\bigl{\{}\hat{L}(X_{i})\leq\tilde{T}_{i}\bigr{\}},$ and $\frac{1}{|\mathcal{I}_{\mathrm{test}}|}\sum_{i\in\mathcal{I}_{\mathrm{test}}}\mathbb{I}\bigl{\{}\min(\hat{L}(X_{i}),2400)\leq\tilde{T}_{i}\bigr{\}}$, respectively. This evaluation protocol offers a fair comparison under relatively limited computations. To measure the variability introduced by the inherently stochastic calibration procedures, we use the same fixed train/validation/calibration/test split for each run, as in our synthetic experiments.
Results. Figure 2 shows the results across $J=5$ runs. Observe how the Uncalibrated baseline overcovers the time-to-unsafe-sampling, resulting in overly conservative LPBs. The Fixed method, on the other hand, yields invalid LPBs with high variance at low budget levels. In contrast, our Optimized method consistently attains near-nominal coverage rates with small variance. Observe how the LPBs obtained by the Optimized method increase as the budget increases—eventually surpassing the Uncalibrated baseline and coinciding with the Fixed method. Overall, this experiment shows that the Optimized method produces LPBs that are both valid across all tested budgets, and become more informative as the budget grows.
@@
Figure 2: Results of experiments on the RealToxicityPrompts dataset as a function of average budget per prompt $B/|{\mathcal{I}_{2}}|$. Left: Empirical coverage rate (target 90%; gray dashed line). The actual coverage of the Optimized method is in solid red line, while the upper and lower bounds on the coverage for the Uncalibrated and Fixed methods correspond to the dotted lines. Right: Mean LPB. Shaded regions represent the standard deviation, computed over $5$ runs. Higher is better.
@@ 7 Discussion
We presented a novel approach to construct a calibrated LPB for time-to-unsafe-sampling of a given prompt, casting this task as a survival analysis problem. Key to our method is the design of a censoring mechanism that utilizes a given budget constraint on the number of LLM generations.
One limitation of our approach is that the coverage guarantee is marginal and does not hold conditionally on a specific subgroup. Consequently, one cannot isolate prompts with especially low (or high) LPBs and assume that those bounds retain their nominal coverage when conditioned on this selection. In future work, we plan to extend our approach to this conditional inference setting, leveraging the methods of [51, 23] as a foundation for achieving valid selection-conditional coverage.
Another limitation is that our calibration assumes the samples are drawn i.i.d. This assumption might not hold in adaptive or adversarial settings, such as jailbreak attacks [38, 49]. Even in benign real-world deployments, the prompt distribution can shift, potentially undermining the validity of our method. In future research, we aim to propose continual or adaptive recalibration mechanisms that adjust to evolving user behavior and emergent risks, possibly by drawing on ideas from [66, 2, 12].

### v4

@@ 5.4 Optimized mode: our flagship method
Having bounded $\gamma$, our next objective is to minimize the average weight $\bar{w}=\frac{1}{|{\mathcal{I}_{2}}|}\sum_{i\in\mathcal{I}_{2}}\frac{1}{\pi_{i}}$ in order to reduce the variance of $\hat{\alpha}(\tau)$. Recall the budget constraint in (2). For $C_{i}:=\text{Ber}(\pi_{i})\cdot\hat{f}_{\tau_{\text{prior}}}(X_{i})$, we get $\mathbb{E}[\sum_{i\in\mathcal{I}_{2}}C_{i}|\{X_{i}\}_{i\in{\mathcal{I}_{2}}}]=\sum_{i\in\mathcal{I}_{2}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\,\pi_{i}$. In turn, we can minimize $\bar{w}$ subject to the budget constraint by solving the following convex optimization problem:
$\pi^{*}=\underset{{\pi\in[0,1]^{|{\mathcal{I}_{2}}|}}}{\rm{argmin}}\;\frac{1}{|{\mathcal{I}_{2}}|}\sum_{i\in\mathcal{I}_{2}}\frac{1}{\pi_{i}}\ \ \text{s.t.}\ \ \sum_{i\in\mathcal{I}_{2}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\,\pi_{i}\!\leq\!B.$ (10)
@@
In Appendix B, we introduce Algorithm 4, which efficiently solves the above optimization problem. Proposition B.2 ensures this algorithm returns a unique and strictly positive solution $\pi^{*}$ that satisfies the budget constraint (2). Importantly, because the solution $\pi^{*}$ depends on the entire set of calibration prompts $\{X_{i}\}_{i\in{\mathcal{I}_{2}}}$, the resulting weights inherit this dependence. To make this explicit, we denote the weight associated with the $i$-th example as $w(\{X_{j}\}_{j\in\mathcal{I}_{2}},i)=1/\pi_{i}^{*}$. We remark that our coverage guarantee from Theorem 4.1 holds also for this formulation of the weights; see Appendix D.6 for details. Furthermore, these weights are also guaranteed to be upper-bounded by the same expression for $\gamma$ as in the Trimmed method.
@@
Suppose that $\max_{i\in{\mathcal{I}_{2}}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\leq M$, then, the weights
$w(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)$ induced by solving (10) are upper bounded by $\gamma=\max\left({|{\mathcal{I}_{2}}|\cdot M}/{B},1\right).$
All proofs are deferred to Appendix D. The above proposition shows us that the solution $\pi^{*}$ inherits the advantages of the Trimmed method and further improves upon it by achieving lower variance, as it minimizes $\bar{w}$. The Optimized calibration scheme is outlined in Algorithm 3.
@@ 6 Experiments
We evaluate our methods on both synthetic (Section 6.1) and real (Section 6.2) datasets. For a full description of how we train models to estimate the conditional quantiles of $T$ given $X$, along with additional experiments, implementation details, and runtime considerations for our calibration methods, see Appendix E. Since the problem setting introduced in this work is novel, there are no established baseline methods against which we can directly compare our performance. Therefore, we consider the vanilla, uncalibrated quantile regression model, as well as the calibrated Naive, Basic, and Trimmed methods as competitive baselines for our Optimized method. The following experiments indicate that our Optimized budget allocation leads to more informative LPBs that attain a coverage rate closer to the desired level compared to these baseline alternatives.
@@ 6.1 Synthetic data experiments
Figure 1: Synthetic experiments. Left: Coverage (target 90%). Center: Mean number of samplings generated per prompt. Right: Mean LPB. Shaded regions represent the standard deviation over 20 runs.
@@
We generate a dataset that simulates a high-risk setting with predominantly prompts that can yield unsafe outcomes. To this end, we generate n=100,000n=100{,}000 pairs $(X_{i},p_{i})$, where $p_{i}$ is the true probability of sampling an unsafe $Y_{i}$ for $X_{i}\in\mathbb{R}^{d}$, with $d=10$. Specifically, $X_{i}\mid p_{i},$ follows a normal distribution, defined in Appendix E.1. The unsafe probability $p$ is chosen to simulate a dataset of “suspicious” prompts, where 90% of which have a high probability for unsafe generation , with the remaining 10% of prompts having a low probability for such an event. Then, we draw $T_{i}\sim\mathrm{Geom}(p_{i})$. We randomly partition the data into training (45%), calibration (45%), and test (10%) sets. For each example in the training set, we generate $500$ outputs. We then use this training set to fit the quantile estimators $\hat{q}_{\tau}(x)$.
@@
We compare (1) the Uncalibrated baseline, the raw quantile estimate $\hat{q}_{\tau}(X)$ at quantile level $\tau=\alpha$; (2) the Naive calibration method, which is a slight variant of the method from Section 5.1. The original Naive approach tended to be unstable, often yielding trivial LPBs, since $\hat{q}_{\tau}(X_{i})\leq C_{i}$ holds for very few points at high quantile levels and these samples do not contribute to the miscoverage estimator in (6). To mitigate this, we restrict the $\tau$’s search space in Algorithm 3 to $\mathcal{T}\cap[0,{\tau_{\textup{prior}}}]$, as in the adaptive budget allocation methods. We also include (3) the Basic method from Section 5.2; (4) its Trimmed version from Section 5.3; and (5) our flagship Optimized approach from Section 5.4.
@@
Recall that the calibration procedures are inherently random due to the sampling of $C_{i}$ and the random generations obtained by the model $\mathcal{G}(x)$. Ideally, this randomness should have a small effect on our bounds. To quantify this randomness, we fix the data split, and run each method for $J=20$ random draws of censoring and time-to-unsafe-sampling times. We index each run by $j$, and denote the resulting LPB by $\hat{q}_{\hat{\tau}}^{(j)}(X_{i})$ and compute:
$\mathrm{AvgCoverage}^{(j)}=\frac{1}{|\mathcal{I}_{\mathrm{test}}|}\sum_{i\in\mathcal{I}_{\mathrm{test}}}\mathbb{P}\{T_{i}\leq$ $\hat{q}_{\hat{\tau}}^{(j)}(X_{i})|X_{i}\},$
$\mathrm{AvgLPB}^{(j)}=\frac{1}{|\mathcal{I}_{\mathrm{test}}|}\sum_{i\in\mathcal{I}_{\mathrm{test}}}\hat{q}_{\hat{\tau}}^{(j)}(X_{i}),\mathrm{AvgBudget}^{(j)}$ $=\frac{1}{|\mathcal{I}_{\mathrm{test}}|}\sum_{i\in\mathcal{I}_{\mathrm{test}}}C_{i}^{(j)}.$
@@
Above, $\mathcal{I}_{\mathrm{test}}$ are the indices of the test points. We report the average and standard deviation of the above quantities over $J$ runs.
@@
Results. We study the effect of the average budget per prompt $B/|{\mathcal{I}_{2}}|$ on the coverage and mean LPB obtained by each method. Following Figure 1, we can see that the Uncalibrated method does not attain valid coverage. The Naive method severely undercovers $T$ for small budgets, and overcovers $T$ for larger budgets. The Basic method attains more stable coverage around the desired $90\%$ level under medium-to-high budget, but exhibits high variance. By contrast, the Trimmed method attains nearly $90\%$ coverage across all budget constraints. Our flagship Optimized calibration method attains tight 90% coverage with minimal variance compared to all other methods. The middle panel highlights that both the Naive and Optimized approaches fully utilize the budget; however, the latter is tightly regulated around the desired coverage—better utilizing the sampling budget. The right panel demonstrates that the Trimmed and Optimized methods produce stable LPBs with better statistical efficiency as the average budget per prompt increases. The increased budget allows us to set higher quantile trim values $M$ while enforcing low $\gamma$. As the budget increases, the two adaptive methods converge to the performance of the Basic technique. However, this technique yields higher LPBs with a higher coverage variability. Additional experiments, including oracle models, varying maximal weights, $\alpha$ levels, and calibration/test splits, are in Appendix E.1.
@@ 6.2 Real data experiments
Figure 2: RealToxicityPrompts dataset experiment. Left: Empirical coverage rate (target 90%). The true coverage of the Optimized method is in a solid red line, while the upper and lower bounds on the coverage of the Uncalibrated and Naive methods correspond to the dotted lines. Right: Mean LPB. Shaded regions represent the standard deviation, computed over $5$ runs. Higher is better.
We now evaluate the effectiveness of our Optimized calibration method on real-world data. For this purpose, we use the RealToxicityPrompts dataset (Gehman et al., 2020), which contains 99,442 sentences designed to measure model toxicity. Each sentence is divided into two halves: the first half serves as the input prompt $X_{i}$ to a language model, which must then complete the sentence. In this experiment, we generate about 30 million model responses. Because this scale is computationally demanding, we employ a relatively small model, Llama 3.2 1B (Meta AI, 2024, Grattafiori et al., 2024). Furthermore, as we do not have access to human annotations, we rely on the Detoxify-original classifier (Hanu & team, 2020) to label each completion as safe (non-toxic) or unsafe (toxic), using a toxicity score threshold of 0.5.
We randomly split the dataset into training (50%), validation (10%), calibration (20%), and test (20%) sets. On the training set, we sample $N=500$ responses per prompt and fit a quantile estimator $\hat{q}_{\tau}(x)$ as detailed in Appendix E. We employ the three calibration methods described in Section 6.1: (1) the Uncalibrated baseline, (2) the Naive method, and (3) the Optimized method, all aiming to attain 90% coverage rate.
In all experiments, we report the empirical mean of the LPBs over the test set. We also estimate the empirical miscoverage by drawing $\min\bigl(\hat{L}(X_{i}),2400\bigr)$ samples for each prompt. We cap the maximal LLM generations on the test set at 2,400 samplings per prompt due to computational constraints. Since the Optimized method trims the quantiles at $M<2400$ across all evaluated budget levels, by drawing exactly $\hat{L}(X_{i})$ samples (with $\hat{L}(X_{i})\leq M$) we can compute the miscoverage rate directly. Since $\hat{L}(X_{i})$ is unbounded for the Uncalibrated and Naive methods, we can only report a lower bound on the empirical coverage rate using the 2,400 generations; see also (Gui et al., 2024). In turn, the empirical lower and upper bounds are given by $\frac{1}{|\mathcal{I}_{\mathrm{test}}|}\sum_{i\in\mathcal{I}_{\mathrm{test}}}\mathbb{I}\bigl\{\hat{L}(X_{i})\leq\tilde{T}_{i}\bigr\},$ and $\frac{1}{|\mathcal{I}_{\mathrm{test}}|}\sum_{i\in\mathcal{I}_{\mathrm{test}}}\mathbb{I}\bigl\{\min(\hat{L}(X_{i}),2400)\leq\tilde{T}_{i}\bigr\}$, respectively. This evaluation protocol offers a fair comparison under relatively limited computations. To measure the variability introduced by the inherently stochastic calibration procedures, we use the same fixed data split for each run, as in our synthetic experiments.
Results. Figure 2 presents the results across $J=5$ runs. The Uncalibrated baseline overcovers the time-to-unsafe-sampling, resulting in overly conservative LPBs. The Naive method, on the other hand, generates invalid LPBs with high variance at small budget levels. In contrast, our Optimized method consistently attains near‐nominal coverage rates with low variance. Notably, the LPBs produced by the Optimized method increase as the budget increases, approaching the values of the Naive and surpassing those of the Uncalibrated method. Overall, this experiment shows that the Optimized method produces LPBs that are both valid across all tested budgets, and become more informative as the budget grows.
@@ 7 Discussion
We introduced a novel approach to construct a calibrated LPB for time-to-unsafe-sampling of a given prompt, casting this task as a survival analysis problem. Key to our method is the design of an optimized censoring mechanism that utilizes a given budget constraint on the number of LLM generations. One limitation of our approach is that the coverage guarantee is marginal and does not hold conditionally on a specific subgroup. Consequently, one cannot group prompts with especially low (or high) LPBs and assume that those bounds retain their nominal coverage when conditioned on this selection. In future work, we plan to extend our approach to this conditional inference setting, leveraging the methods by Sesia & Svetnik (2025), Jin & Ren (2025) as a foundation for achieving valid group-conditional coverage. Another limitation is that our calibration assumes the samples are drawn i.i.d. Even in benign real-world setups, the prompt distribution can shift, potentially undermining the validity of our method. In future research, we aim to propose continual or adaptive recalibration mechanisms that adjust to evolving user behavior and emergent risks, possibly by drawing on ideas from (Gibbs & Candès, 2024).